OpenStudy (anonymous):

what is 5 1/4 divided by 1 5/9??

OpenStudy (anonymous):

first convert the mixed number into improper fractions

OpenStudy (ribhu):

convert in proper fractions.

OpenStudy (alexandervonhumboldt2):

to convert to proper fraction multiply number by denomiantor and then add numenator to it

OpenStudy (ribhu):

multiply the 1st fraction with the reciprocal of 2nd fraction.

OpenStudy (anonymous):

how?

OpenStudy (ribhu):

5 1/4 = 21/4; 1 5/9 = 14/9

OpenStudy (anonymous):

an then do i figure out how many times 9 goes into 14 and thats my answer??

OpenStudy (anonymous):